Find the midpoint between the points (7, -7) and (-3, 3).
Answer by checkley71(8403)

FOR THE X MID POINT 7+3=10/2=5 OR 7-5=2 FOR THE X MID POINT.
FOR THE Y MID POINT 3+7=10/2=5 OR 3-5=-2 FOR THE Y MID POINT.
ANSWER (2,-2).
PROOF IS IN THE GRAPHING OF THE 2 POINTS, DRAWING A LINE BETWEEN THEM & DIVIDING THE LINE IN HALF.